# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ Abstract
Multimodal contrastive representation learning methods have proven successful across a range of domains, partly due to their ability to generate meaningful shared representations of complex phenomena. To enhance the depth of analysis and understanding of these acquired representations, we introduce a unified causal model specifically designed for multimodal data. By examining this model, we show that multimodal contrastive representation learning excels at identifying latent coupled variables within the proposed unified model, up to linear or permutation transformations resulting from different assumptions. Our findings illuminate the potential of pre-trained multimodal models, e.g., CLIP, in learning disentangled representations through a surprisingly simple yet highly effective tool: linear independent component analysis. Experiments demonstrate the robustness of our findings, even when the assumptions are violated, and validate the effectiveness of the proposed method in learning disentangled representations.
Machine Learning, ICML
@@ 1 Introduction
As humans, we often encounter situations where being aware of a signal from one modality helps in the interpretation of another. For instance, observing someone’s facial expressions while they speak facilitates comprehension beyond merely listening to the audio. Similarly, in machine learning, there are often instances where data from one modality can enhance the interpretation of another. This becomes particularly valuable in situations where obtaining explicit training labels is challenging or resource-intensive (Quattoni et al., 2007; Shi et al., 2019). One promising strategy in this context is to use data from one modality, e.g., text data, as a supervision signal in the interpretation of another, e.g., image data (Mori et al., 1999; Wang et al., 2009; Ramanathan et al., 2013; He & Peng, 2017; Radford et al., 2021). The primary approach for achieving this is known as multimodal contrastive representation learning, which focuses on optimizing a symmetric contrastive loss (Zhang et al., 2022; Radford et al., 2021), e.g., a symmetric adaptation of the standard contrastive loss (Wu et al., 2018; Tian et al., 2020; He et al., 2020; Chen et al., 2020). The learned representations, guided by the symmetric contrastive loss, have been applied in a variety of applications, including zero/few-shot learning (Radford et al., 2021; Zhou et al., 2022a), domain generalization (Zhou et al., 2022a, b), and robustness to adversarial examples (Ban & Dong, 2022). Despite its proven success in practice, understanding the effectiveness of learned representations presents a non-trivial challenge.
Previous studies investigated the effectiveness of standard contrastive learning (on a single modality) by exploring the connection between the learned representations and the latent variables in latent causal generative models (Zimmermann et al., 2021; Von Kügelgen et al., 2021). This exploration offers a vital causal viewpoint, highlighting the use of independent causal mechanisms (Schölkopf et al., 2021; Rojas-Carulla et al., 2018) to understand the generalization ability of these learned representations across various downstream applications. However, these studies are predominantly restricted to single-modality scenarios, typically images. The nature of the relationship between the representations learned by the symmetric contrastive learning in multimodal settings and the latent variables within latent causal generative models remains under-explored. Further related works are detailed in Appendix A.1.
The existence of the close connection between the representations learned in multimodal contrastive learning and the latent variables in latent causal generative models can be revealed from the symmetric contrastive loss and recovering the latent variables. Specifically, the goals of multimodal contrastive learning, i.e., aligning real pairs and separating incorrect ones (Zhang et al., 2022; Radford et al., 2021), mirror the essential objectives in solving inverse problems. In this perspective, aligning real pairs in multimodal contrastive learning is closely related to matching a type of prior information on latent variables while separating the incorrect pairings, aiming for capturing the full range and complexity of the latent variables. Accordingly, we demonstrate the connection between the two paradigms formally in the following.
We introduce a novel latent partial causal model to formulate multimodal contrastive learning, with the reflection of their natural connection. It formulates the causal generative process of the paired observations and the latent variables capturing the modality-specific and shared patterns. Specifically, we propose to capture the cross-modality shared patterns relying on a coupling relation between a couple of separate latent variables. Unlike previous works employing a single latent variable to model the shared patterns (Daunhawer et al., 2023; Gresele et al., 2020; Yao et al., 2023), the proposed method uses coupling relation and separate latent variables (i.e., latent coupled variables) to capture the flexible cross-modality matching relationship in practical multimodal contrastive learning (e.g., CLIP) more generally. Apart from the flexibility and generalization, furthermore, benefiting from the modeling with separate and coupled latent variables, we show that all latent coupled variables in the proposed causal model can be identified up to linear or permutation transformations, subject to specific statistical assumptions. This finding sets our work apart from prior studies focused on partial and block identifiability (Daunhawer et al., 2023; Yao et al., 2023), e.g., identifying part of latent coupled variables up to an invertible nonlinear mapping, thus offering a more comprehensive and fine-grained understanding for representations learned by multimodal contrastive representation learning.
The theoretical insights above reveal a remarkable facet of multimodal contrastive learning, showcasing their inherent and potent capacity to learn disentangled representations, under specific conditions. This remarkable ability can be brought to light for multimodal contrastive learning and the pre-trained models, such as CLIP (Radford et al., 2021), through the application of a straightforward yet highly effective tool, linear independent component analysis (ICA) (Hyvärinen et al., 2001). This provides the first step towards a guarantee for the disentanglement ability of multimodal contrastive representation learning. We validate our theoretical findings under ideal assumptions, and illustrate the robustness of our findings even when the assumptions are partially violated. We conduct experiments on real datasets, demonstrating the disentanglement ability of pre-trained CLIP when combined with the Fast ICA method (Hyvarinen, 1999), in terms of learning disentangled representations, few-shot learning and domain generalization.
2 Motivation and The Proposed Model
In this section, we begin by providing a basic motivation for the connection between representations learned by multimodal contrastive representation learning and the latent variables in latent causal generative models in Section 2.1. Building upon this motivation, we then introduce a novel latent partial causal model and clarify how this model differs from previous works, specifically those by (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2023; Gresele et al., 2020; Yao et al., 2023), in Section 2.2.
2.1 Motivation: Relation with Inverse Problem
Multimodal contrastive representation learning leverages latent shared patterns in multimodal data, by utilizing information from one modality to supervise another. Recent developments in this field show that representations learned through a symmetric contrastive loss are effective in various downstream tasks, such as transfer learning, and are robust to shifts in natural distributions (Radford et al., 2021). The contrastive loss function is designed to maximize the similarity in the embedding space between modalities for real pairs, while minimizing the similarity for incorrect pairs. Formally, the optimization objective is as follows:
$\mathcal{L}=$ $-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$ (1)
$-\frac{1}{N}\sum_{i=1}^{N}\log\frac{{e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}}{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}},$
@@ 2.1 Motivation: Relation with Inverse Problem
where $d$ denotes a distance metric, e.g., cosine similarity on hypersphere or L1 norm on convex bodies, $\tau$ is a learnable temperature parameter, $N$ denotes the sample size, which means that we have $N$ positive pairs and $N^{2}-N$ negative pairs, $\mathbf{f}_{x}$ denote the encoder on one modality $\mathbf{x}$, e.g., image, similarly, $\mathbf{f}_{t}$ denote the encoder on another $\mathbf{t}$, e.g., text.
To further understand the symmetric contrastive loss, we investigate its asymptotics as follows:
@@
For fixed $\tau>0$, as the sample size $N\rightarrow\infty$, the (normalized) symmetric contrastive loss converges to
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$ (2)
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$
Proof sketch
The proof can be accomplished by adapting the proof methodology utilized in Theorem 1 of Wang & Isola (2020) for a single modality to the context of multiple modalities. It primarily relies on the application of the strong law of large numbers and the continuous mapping theorem. Refer to Appendix A.2 for details.
Insight
Our primary insight is that the loss function Eq. (2) is intricately linked to two fundamental elements crucial for solving inverse problems, e.g., with nonlinear ICA: 1) Matching prior information on latent variables, i.e., the solution space is constrained by prior distribution, helping to mitigate issues of non-uniqueness in inverse problems. 2) Recovering all information in latent space, e.g., ensuring that the inverse problem solution captures the full range and complexity of the latent variables.
Relation with Matching Prior Information
To clarify this point, let us focus on the first term at the right-hand side of Eq. (2). In a multi-model setting, one modality plays a crucial role as a supervised signal for another modality. This implies that one modality can serve as potential prior knowledge. Let us consider the modality of text as this prior knowledge. By minimizing the first term, which reduces the distance between features obtained by encoders for real pairs, we essentially ensure the features generated by one encoder, e.g., on the image modality, closely approximate the prior knowledge provided by the text. This closely aligns with a key element essential for solving inverse problems: matching prior information. For instance, when recovering latent variables from observational data, we expect the recovered variables to closely approximate the prior knowledge derived from the true latent variables.
Relation with Recovering All Information
To elucidate this point, let us focus on the last two terms on the right-hand side of Eq. (2). Essentially, these two terms are approximately equal to optimizing the following (Proof can be found in Appendix A.3):
@@
Here, $H(\cdot,\cdot)$ represents the cross entropy. When both $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ transform $\mathbf{x}$ and $\mathbf{t}$ into uniformly distributed random variables, respectively, Equation (3) reaches its optimal solution. This uniform distribution underscores our goal of finding functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ that maximize information preservation. This objective aligns with a crucial aspect of solving inverse problems, which is to recover all latent space information.
Expanding upon the insight presented, we posit that multimodal contrastive representation learning has the potential to identify certain latent variables. The pertinent question is what latent variables in a latent causal generative model are identified. Subsequent sections will introduce a novel latent causal generative model to elucidate this connection.
2.2 The Proposed Latent Partial Causal Model
Figure 1: Illustration of the proposed latent partial causal model: The latent space is partitioned into $(\mathbf{m}_{x},\mathbf{z}_{x})$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$, where $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$ represent modality-specific latent variables. An undirected edge between $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ is employed to model latent shared patterns. The observations $\mathbf{x}$ (e.g., images) and $\mathbf{t}$ (e.g., text) are generated by two distinct generative processes, $\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})$ and $\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$, respectively.
Figure 1 illustrates the proposed latent partial causal model. In this model, the whole latent space is partitioned into two parts, each representing a different modality, e.g., image and text. More specifically, to model latent shared patterns, an undirected edge is established between $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ to allow subtle perturbation, which will be formulated by conditional distributions in the next section. The rationale behind this modeling approach is grounded in the recognition that real-world multimodal data is often complex, noisy, and multifaceted. On one hand, the assertion ‘a picture is worth a thousand words’ is well-supported in literature (Gropper, 1963; Hum et al., 2011), emphasizing the rich detail and information that images can convey compared to text. Conversely, this notion is not universally applicable as argued by Reinert (1976), who suggests that sometimes textual information can be more informative than visual data. This perspective is further echoed by Fidler et al. (2013) in their assertion that ‘a sentence is worth a thousand pixels,’ highlighting the potential of text in conveying complex ideas succinctly.
In addition, we introduce $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$ modality-specific latent variables, each tailored to capture the unique characteristics of their respective domains. For example, $\mathbf{m}_{x}$ could encode information focusing on aspects like the presence of background noise or other visual artifacts that contribute to the overall composition of an image. On the other hand, $\mathbf{m}_{t}$ could encode information about sentence structure or linguistic patterns that are characteristic of the grammar in textual content. Finally, the observations are associated with two distinct generative processes. Specifically, $\mathbf{x}$ (e.g., images) are generated through the process $\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})$, while observation $\mathbf{t}$ (e.g., text) come into existence through the generative process $\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$.
Differences with previous works
The proposed model distinguishes itself from previous works such as (Zimmermann et al., 2021; Von Kügelgen et al., 2021), which leverage latent causal models to elucidate contrastive learning. Prior efforts predominantly concentrate on single-modal settings, while our proposed model extends this perspective to the more intricate domain of multi-modal settings. Loosely speaking, the proposed model can be considered a generalization of previous works. This generalization is apparent as the proposed model seamlessly reduces to a single-modal setting when the mixing functions from latent space to observed space are enforced to be identical, and specific variables are omitted. Furthermore, our model differs from previous works specifically designed for multi-modal settings (Daunhawer et al., 2023; Gresele et al., 2020; Yao et al., 2023), a distinction that becomes evident across two key dimensions: 1) Modeling shared patterns: the proposed model captures shared patterns across modalities allowing for the existence of a coupling relation in multimodal data. In contrast, previous works often achieve this by introducing the same variables. Notably, our modeling approach using the coupling relation is more general, as it can be reduced to using the same variables by enforcing an identical mapping on the coupling relation. 2) Identifiability results: The identifiability results obtained from the proposed model diverge from those found in previous works (Daunhawer et al., 2023; Yao et al., 2023), both in terms of breadth and depth of identifiability, due to the introduction of the coupling relation (will be demonstrated in Section 3). a) (Breadth of Identifiability) Unlike earlier models that often achieve only partial identifiability of latent coupled variables, e.g., latent content variables but not latent style variables (Daunhawer et al., 2023), our model extends this scope to ensure complete identifiability of coupled variables. This comprehensive identifiability result allows for a more complete understanding and representation of the data. b) (Depth of Identifiability) In terms of depth, our model provides a more nuanced and detailed analysis by identifying coupled variables up to linear or permutation transformations. This level of precision offers an enhancement over the block identifiability approach typically seen in previous studies (Daunhawer et al., 2023; Yao et al., 2023), allowing for learning disentangled representations. We will further clarify this in section 4.
3 Identifiability Analysis
In this section, we conduct an identifiability analysis for the proposed latent partial causal model illustrated in Figure 1. Our analysis specifically focuses on two distinct types of latent spaces: the hypersphere and convex bodies, which are explored under certain defined assumptions.

### v2

# Beyond DAGs: A Latent Partial Causal Model for Multimodal Learning

@@ Abstract
Directed acyclic graphs (DAGs) are fundamental graph structures in causal modeling, but identifying the desired DAG from observational data often requires strong assumptions that may not hold in real-world scenarios, especially for latent causal models and complex multimodal data. This raises the question of whether we can relax or bypass the DAG assumption while maintaining practical utility. In this work, we propose a novel latent partial causal model for multimodal data, featuring two latent coupled variables, connected by an undirected edge, to represent the transfer of knowledge across modalities. Under specific statistical assumptions, we establish an identifiability result, demonstrating that representations learned by multimodal contrastive learning correspond to the latent coupled variables up to a trivial transformation. This result deepens our understanding of the why multimodal contrastive learning works, highlights its potential for disentanglement, and expands the utility of pre-trained models like CLIP. Synthetic experiments confirm the robustness of our findings, even when the assumptions are partially violated. Most importantly, experiments on a pre-trained CLIP model embodies disentangled representations, enabling few-shot learning and improving domain generalization across diverse real-world datasets. Together, these contributions push the boundaries of multimodal contrastive learning, both theoretically and, crucially, in practical applications.
[…]
@@ 1 Introduction
The adoption of a directed acyclic graph (DAG) as the underlying graph structure is a foundational assumption in most causal modeling methods [45, 55]. Identifying the DAG typically requires additional assumptions that enforce the asymmetry between nodes, allowing for the distinction of causes and effects [13]. In causal discovery within the observed space, these assumptions often include imposing constraints on the function class, such as linear non-Gaussian [53] or additive noise models [21, 46]. However, due to the complex and uncertain nature of real-world data, justifying these restrictive function classes is often impractical.
This challenge becomes even more pronounced in latent spaces, such as in causal representation learning (CRL) [51], where the relationships among observational data are assumed to arise from causal interactions between high-level latent causal variables. Most existing methods for identifying latent causal variables rely on the assumption that data resulting from interventions on all latent variables are available [36, 4, 5, 57, 1, 52, 37]. However, acquiring such interventional data can be difficult, and in many cases, it may be impossible in real-world scenarios.
This challenge is further amplified in multimodal settings, as the causal relationships across different modalities often involve intricate dependencies that may not conform to a single global DAG structure. This raises a fundamental question: can we develop a more general latent causal modeling framework for multimodal settings that relaxes or circumvents DAG structure assumption while maintaining practical utility in real-world scenarios?
We introduce a novel latent partial causal generative model in this work, to characterize the multimodal data generative process, as depicted in Figure 1. Unlike traditional DAG assumptions, our framework introduces latent coupled variables linked by an undirected edge to represent the transfer of knowledge across modalities, enabling flexible modeling of complex dependencies. Additionally, the model incorporates modality-specific latent variables to capture distinct information within each modality and employs dedicated mappings from the latent space to the observed space, facilitating data generation tailored to each modality.
Building on the proposed latent partial causal generative model, we examine its relationship with multimodal contrastive learning, a widely-adopted inference model [64, 47]. Through identifiability analysis, we establish a theoretical link between the generative and inference models, showing that the representations learned by multimodal contrastive learning correspond to the latent coupled variables in the generative model. This correspondence holds up to a trivial linear transformation in hypersphere space and a permutation transformation in convex bodies. These results provide a theoretical foundation for the success of multimodal contrastive learning.
Furthermore, our linear identifiability results in hypersphere space and permutation identifiability in convex bodies provide strong theoretical justification for the disentanglement capabilities of multimodal contrastive learning, under the assumption that each latent coupled variable is mutually independent. This offers a surprising insight into leveraging the disentanglement potential of CLIP-like models [47], trained by multimodal contrastive learning, to address tasks that rely on disentangled representations. For instance, this potential can be utilized to enable the learning of disentangled representations, enhance performance in few-shot learning and domain generalization.
In summary, our contributions are as follows:
  • We propose a novel latent partial causal model for the generative process, specifically designed for multimodal data. Instead of relying on DAGs assumption, our model introduces latent coupled variables, connected by undirected edges, to effectively capture transferable knowledge across modalities.
  • Our identifiability analysis demonstrates that the representations learned by multimodal contrastive learning align with latent coupled variables in the generative model. This insight highlights that the success of multimodal contrastive learning lies in its capacity to identify these latent variables within generative models.
  • Furthermore, our identifiability results establish the theoretical basis for the component-wise disentanglement potential of multimodal contrastive learning. This pushes the boundaries of how pre-trained models, such as CLIP, can be leveraged. To the best of our knowledge, this is the first work to provide guarantees for the component-wise disentanglement potential.
  • We validate our theoretical findings under ideal conditions and demonstrate their robustness even when assumptions are partially violated. Extensive experiments on pre-trained CLIP model across various tasks—such as few-shot learning, domain generalization, and disentangled representation learning—on over 16 real-world datasets substantiate the practical effectiveness of latent coupled models. We emphasize that, unlike prior works on identifiability analysis for multimodal learning [8, 62, 14], which are limited to simulation experiments, this work empirically validates identifiability results using real datasets and models, thereby demonstrating the robustness and practical significance of our theoretical findings.
In conclusion, the proposed latent partial causal model supports the success of multimodal contrastive learning and reveals its potential for disentanglement. Most importantly, experiments on pre-trained CLIP across various tasks using extensive real-world data demonstrate the practical utility of our theoretical results. Collectively, these findings relax the conventional reliance on DAG assumptions in multimodal contrastive learning, while maintaining practical utility in real-world scenarios.
2 Generative Model: The Latent Partial Causal Model
Figure 1: The proposed latent partial causal model. $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ are latent coupled variables, and $\mathbf{m}_{x}$, $\mathbf{m}_{t}$ are modality-specific.
In this section, we introduce a novel latent partial causal model designed to represent the generative process of multimodal data. Unlike traditional DAG assumptions, our model incorporates an undirected edge between two latent variables, enabling the capture of transferable knowledge across modalities, as depicted in Figure 1. This non-DAG structure offers substantial flexibility, allowing the representation of multiple DAG assumptions as special cases, as illustrated in Figure 2. By doing so, our model facilitates the extraction of transferable latent knowledge that underpins the fundamental principles behind various DAG assumptions, thereby enabling a more flexible and generalized approach to multimodal causal modeling.
The Proposed Latent Partial Causal Models
Figure 2: Illustrative DAGs behind the proposed model: Left: A latent confounder influences both $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. Middle: $\mathbf{z}_{t}$ influences $\mathbf{z}_{x}$ through an intermediate mediator $\mathbf{b}$, serving as a bottleneck for transferable knowledge. Right: A symmetric inverse relationship where $\mathbf{z}_{x}$ influences $\mathbf{z}_{t}$ via $\mathbf{b}$.
Figure 1 illustrates the proposed latent partial causal model, designed to represent the generative process for multimodal data. The latent space is partitioned into two components, each corresponding to a specific modality, such as image and text. To capture unique characteristics within each domain, the model incorporates modality-specific latent variables, $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$ tailored. For instance, $\mathbf{m}_{x}$ may represent domain-specific aspects such as background noise or visual artifacts, while $\mathbf{m}_{t}$ could encode linguistic features such as sentence structure or grammatical patterns. Further, the observations are generated through distinct processes that link the latent variables to the observed data. Specifically, images ($\mathbf{x}$) are generated by the function $\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})$, while text ($\mathbf{t}$) is produced by $\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$.
In addition, to capture transferable knowledge between these modalities, the model introduces an undirected edge between the latent coupled variables, $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. The rationale behind this modeling approach is grounded in the intricate dependencies between modalities. For instance, the adage “a picture is worth a thousand words” highlights the richness and detail of visual data, as supported by [15, 23]. However, this perspective is not universally applicable, as [49] argues that textual information can often convey more precise meanings. Similarly, [11] reinforces the complementary nature of text, asserting that “a sentence is worth a thousand pixels” in its ability to succinctly express complex ideas.
Beyond the rationale outlined above for introducing coupled variables and an undirected edge between them, two key factors justify this design: 1) Modality Gap: In multimodal models like CLIP [35], different data modalities are often embedded at a distance in the shared representation space. This suggests that using two coupled variables to represent transferable knowledge across modalities may be appropriate. 2) The diversity of real-world multimodal data: Real-world multimodal data often entails complex dependencies that a single global structure may fail to capture. More importantly, these dependencies could even be represented by conflicting directed acyclic graphs (DAGs). Figure 2 illustrates such potential DAGs, offering insights into the diverse underlying causal structures.
In the left DAG of Figure 2, the latent confounder $\mathbf{c}$ represents a shared source of variation that influences both $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. This confounder captures a common underlying context or concept connecting the two modalities. For example, if the image and text are related to the topic "sports," $\mathbf{c}$ could encapsulate this shared theme, influencing the generation of both the visual and textual data. The middle DAG depicts a structure where $\mathbf{b}$, representing transferable knowledge. Specifically, $\mathbf{b}$ serves as the bridge, deriving information from the text latent variable $\mathbf{z}_{t}$ and informing the image latent space $\mathbf{z}_{x}$. This scenario aligns with tasks where text serves as a guiding input for image generation, e.g., text-to-image generalization. A classical example is the MNIST dataset. In contrast, the DAG on the right represents an image-guided text generation process, e.g., image captioning. Here, the high-level latent information in the image influences the high-level latent variable in the generated caption. A classical example is the CelebA-Dialog dataset [31].
3 A First Look: The Recovery Potential of Multimodal Contrastive Learning
Given the proposed latent coupled generative model, as depicted by Figure 1, our final goal is to analyze how multimodal contrastive learning framework can recover the true latent coupled variables, up to an sample transformation. Before this, we first provide an intuitive motivation to explain why we believe multimodal contrastive learning can achieve this.
Multimodal contrastive learning leverages a loss function designed to maximize similarity between embeddings of real paired data while minimizing similarity for incorrect pairs. The loss function is defined as [64, 47]:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}-\frac{1}{N}\sum_{i=1}^{N}\log\frac{{e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{j}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}},$ (1)
@@ 3 A First Look: The Recovery Potential of Multimodal Contrastive Learning
where $d$ denotes a distance metric, e.g., cosine similarity on hypersphere or L1 norm on convex bodies, $\tau$ is a learnable temperature parameter, $N$ denotes the sample size, which means that we have $N$ positive pairs and $N^{2}-N$ negative pairs, $\mathbf{f}_{x}$ denote the encoder on one modality $\mathbf{x}$, i.e., image, similarly, $\mathbf{f}_{t}$ denote the encoder on another $\mathbf{t}$, i.e., text.
To understand the multimodal contrastive loss further, we investigate its asymptotics:
@@
For fixed $\tau>0$, as the sample size $N\rightarrow\infty$, the (normalized) multimodal contrastive loss converges to
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=$ $2\operatorname*{\mathbb{E}}\limits_{(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$ (2)
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$
This result extends Theorem 1 from [60] to the multimodal context. Proof is provided in Appendix B.
Insights into Latent Variable Recovery
The loss function in Eq. (2) connects directly to two fundamental principles in latent variable recovery: Prior Matching and Information Preservation. These principles are crucial for methods like nonlinear independent component analysis (ICA) [28], which recover latent independent variables from observed data.
  • Prior Matching: This constrains the solution space using prior knowledge, addressing the non-uniqueness problem that often arises in latent variable recovery.
  • Information Preservation: This ensures that the solution space fully captures the complexity of the latent variables derived from the observed data.
Prior Matching
The first term in Eq. (2) promotes alignment between representations of real data pairs across modalities, enforcing that one modality (i.e., text) acts as a prior signal for the other (i.e., image). Minimizing this term drives cross-modal alignment and incorporates prior knowledge, which is key for recovering latent variables.
Information Preservation
The last two terms in Eq. (2) are closely related to ensuring that the learned representations capture the full complexity of the latent variables. These terms can be approximated by optimizing the following expression (proof in Appendix C):
@@
where $H(\cdot,\cdot)$ denotes cross-entropy. The objective function in Eq. (2) is symmetric between $\mathbf{x}$ and $\mathbf{t}$. Intuitively, if $p(\mathbf{f}_{x}(\mathbf{x}))$ and $p(\mathbf{f}_{t}(\mathbf{t}))$ are not equal, the solution deviates, increasing the objective value and introducing asymmetry in the last two terms. For the optimal solution, the two distributions must align. When $p(\mathbf{f}_{x}(\mathbf{x}))=p(\mathbf{f}_{t}(\mathbf{t}))$, the cross-entropy in Eq. (3) reduces to entropy, and if $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ transform $\mathbf{x}$ and $\mathbf{t}$ into uniformly distributed random variables, Eq. (3) reaches its optimal value. This highlights the importance of finding transformations $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ that preserve information by fully capturing the latent variable structure.
A Novel Unified Perspective on Contrastive Loss
Previous research has primarily focused on contrastive loss in the context of single modalities, emphasizing two main perspectives: 1) alignment-uniformity [60], which is closely related to prior matching, and 2) information preservation [43]. However, these perspectives have largely been treated separately. In this work, we offer a novel insight by combining these two perspectives within the multimodal context for latent variable recovery. This insight motivates our belief that multimodal contrastive representation learning holds significant potential for recovering latent variables.
4 From Potential to Principles: Identifiability Guarantee
Given the initiative analysis in Section 3, which highlights the potential of multimodal contrastive learning for recovering latent variables, we now move forward to rigorous identifiability analysis, which provide theoretical guarantees that multimodal contrastive learning can indeed recover the true latent variables, by parameterizing the proposed latent partial causal model. We examine two distinct types of parameterization in latent spaces, hyperspheres and convex bodies, under specific assumptions, respectively.